# JudgeAgent: Dynamically Evaluate LLMs with Agent-as-Interviewer

### v1

@@ Abstract
Evaluating the capabilities of large language models (LLMs) is an essential step to ensure the successful application of LLMs across various domains. The current evaluation of LLMs is based on a paradigm that involves querying them with predefined question sets and assessing their outputs. This paradigm offers controllable processes and simplicity, but faces challenges such as limited interaction with targets, insufficient difficulty control, and difficulties in verifying the validity of evaluation results, making it hard to precisely determine the knowledge and capability boundaries of target models. To address these challenges, we propose JudgeAgent, a knowledge-target adaptive dynamic evaluation framework based on a new interviewer-style evaluation paradigm. JudgeAgent employs a comprehensive evaluation approach consisting of benchmark grading, interactive extension, and evaluation feedback. It utilizes knowledge-driven data synthesis and target-adaptive difficulty adjustment methods to conduct extended testing, providing accurate and effective evaluation results. We also introduce a novel insight into validating evaluation methods, demonstrating the effectiveness of JudgeAgent and its dynamic evaluation paradigm through extensive experiments.
@@ 1 Introduction
In recent years, large language models (LLMs) have achieved remarkable success across various domains(Xu et al., 2024), from technical fields(Tang et al., 2024; Yuan et al., 2023) to social sciences(Shi, 2024; Wang et al., 2025), serving in multiple roles such as generators(Bubeck et al., 2023; Yu et al., 2024) and evaluators(Liu et al., 2023; Wang et al., 2023). This versatility stems from their massive model scale, extensive training data, and the resulting strong capabilities in comprehension and reasoning, as well as the vast amount of knowledge stored within them. Consequently, assessing the capabilities and weaknesses of LLMs has become an essential step before their deployment.
Current evaluations of LLMs are generally conducted through a paradigm that involves querying the models with predefined question sets and assessing their responses. The evaluation of outputs is typically carried out by comparing them against ground-truth reference answers, through human assessment, or via the emerging approach, LLM-as-a-judge, which uses LLMs to evaluate objects based on predefined rules and criteriaGu et al. (2024). The sources of these question sets mainly fall into two categories. The first is static datasets, which are constructed through manual collection or model-generated questions that undergo quality screening(Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023; Lin et al., 2022; Cobbe et al., 2021; Tang & Yang, 2024). Such datasets offer advantages, including well-controlled question quality, straightforward querying procedures, and thematic consistency. However, their static nature makes them vulnerable to data leakage(Zhou et al., 2023), and easier to be cheated by target models(Zheng et al., 2024). As a result, older datasets may fail to accurately reflect the capabilities of state-of-the-art models. The second category consists of dynamically generated questions, currently also primarily produced by LLMs based on predefined requirements(Bai et al., 2023; Shi et al., 2025; Yu et al., 2024). While this approach largely mitigates the risk of data leakage, issues such as inaccuracies and irrelevance may arise due to inherent limitations in LLMs, including knowledge gaps, biases, and hallucinations(Ji et al., 2023). Furthermore, existing evaluation methods often lack sufficient interaction with the target model and do not incorporate adaptive difficulty control, making it difficult to accurately determine the model’s knowledge and capability boundaries. Moreover, the absence of interpretable feedback hinders the effective application of evaluation results to guide model iteration and optimization, and the validity of the evaluations remains challenging to verify.
To address these challenges, we propose JudgeAgent, a dynamic, knowledge-target adaptive evaluation framework with a new interviewer-style evaluation paradigm. Inspired by how a human interviewer dynamically and deeply probes a candidate’s knowledge and abilities, JudgeAgent employs a novel evaluation paradigm comprising benchmark grading, interactive extension, and evaluation feedback. By leveraging knowledge-driven synthesis of data and adaptively adjusting question difficulty based on the target model, JudgeAgent conducts extensive testing to comprehensively and profoundly assess the target model’s knowledge and ability levels. Furthermore, it provides multi-dimensional and interpretable evaluation feedback to support the target’s optimization and iteration.
@@
  • Interviewer-Style Evaluation Paradigm: We propose a dynamic, strongly interactive evaluation paradigm, which comprehensively and deeply assesses target capabilities through benchmark grading, interactive extension, and evaluation feedback processes.
  • Knowledge-Target Adaptive Evaluation Framework: We introduce JudgeAgent, which dynamically adjusts difficulty based on interactive feedback and employs knowledge-driven synthetic data generation to create questions that calibrate to relevant knowledge boundaries and limits of target capabilities, enhancing evaluation validity.
  • New Insights into Validation of Evaluation: We conduct thorough experiments and analysis to validate the effectiveness of JudgeAgent’s evaluation of various target models across multiple datasets. This work also offers novel insights into how evaluation methodologies themselves can be measured.
@@ 2.1 Static Benchmark-based Evaluation
These methods employ pre-constructed static benchmarks to evaluate LLMs in specific domains or on specific tasks, using formats such as multiple-choice questions, question–answer questions, or prompts that require LLMs to perform designated tasks. For multiple-choice(Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023) or question–answer(Lin et al., 2022; Cobbe et al., 2021; Tang & Yang, 2024) formats, the benchmark provides correct answers as ground truth, and the target’s accuracy in answering the questions is used to measure its capability in the specific domain. For task-execution or open-ended formats, the benchmark typically offers only reference answers, with correctness and plausibility assessed by human evaluators(Chang et al., 2024) or LLM-based evaluators(Wang et al., 2023; Zheng et al., 2023).
The primary advantage of such methods lies in the quality of the evaluation questions, which can be well assured after prior curation. However, their static nature makes them susceptible to data leakage, whereby the target model may have prior exposure to the benchmark and can thus exploit this knowledge, undermining the validity of the evaluation.
@@ 2.2 Dynamic LLM-based Evaluation
These methods leverage LLMs to dynamically generate questions for testing target models, and in some cases, to directly serve as evaluators that provide evaluation results based on the target model’s performance. Compared with static methods, their key advantage lies in the fact that the generated questions are difficult for the target model to anticipate or tailor responses to, thereby producing evaluation results that are more likely to reflect the target’s genuine capabilities rather than its ability to recall memorized items.
For example, Bai et al. (2023) employs an LLM as an examiner, generating questions across diverse domains based on Google Trends categories, and uses follow-up questions to assess both breadth and depth of knowledge. SafetyQuizzer(Shi et al., 2025), on the other hand, retrieves up-to-date news via search engines to construct background texts and then applies retrieval-augmented generation (RAG) to formulate questions grounded in current events, maximizing question timeliness to mitigate cheating risks. These methods primarily focus on the knowledge sources used for dynamic question generation, but often lack extensive interaction and feedback with the target model. KIEval(Yu et al., 2024) emphasizes interaction with the target model by starting with known questions from a dataset and continuously issuing follow-up queries based on the target’s responses, thereby probing the depth of the target’s recall and understanding of knowledge related to the selected question. However, because it remains centered on a single question, its coverage of knowledge breadth is limited.
Moreover, dynamic evaluation methods, including the aforementioned approaches, tend to focus predominantly on the content of the questions themselves, without fully considering the alignment between question difficulty and the target model’s capability range. Evaluations conducted with difficulty levels that deviate from the target’s ability inevitably yield coarse and biased results. To address this limitation, we propose JudgeAgent, which performs authentic and fine-grained capability evaluation through iterative interaction and adaptive difficulty adjustment.
@@ 3 Methodology
Figure 1: The framework of JudgeAgent. The left part represents the interaction between the target model and JudgeAgent. The central part delineates the modular composition and operational mechanics of JudgeAgent. The right part presents the tools upon which JudgeAgent relies.
We propose JudgeAgent, a dynamic evaluation framework based on an ”interviewer-style” evaluation paradigm of benchmark grading, interactive extension, and evaluation feedback. The framework’s workflow comprises three core components: (1) Benchmark Grading: Determining the approximate capability range of the evaluation target through testing on public static datasets. (2) Interactive Extension: Using knowledge-driven synthetic methods, JudgeAgent iteratively adjusts question difficulty based on estimated target capability levels, generates extension questions for evaluation, and dynamically updates capability estimations. (3) Evaluation Feedback: Integrating all test results, JudgeAgent conducts comprehensive capability evaluations, identifies the target’s deficiencies in logical reasoning, semantic understanding, and other aspects of capabilities, and provides actionable optimization suggestions as feedback. The overall framework is illustrated in Figure X, and the detailed process is demonstrated in Algorithm 1 in Appendix B.
@@ 3.1 Benchmark Grading
Analogous to a written test before an interview, JudgeAgent, acting as the interviewer, forms an initial understanding of the target model’s capabilities by evaluating its performance on a publicly available static dataset. This process consists of two stages: Testing and Tiering.
In the Testing stage, for a selected dataset $\mathcal{D}=\{q_{1}\,,\,q_{2}\,$,\,...\}, JudgeAgent partitions a specified number of questions into batches $\mathcal{B}=\{B_{1}\,,\,B_{2}\,$,\,...\} ($B_{i}=\{q_{i1}\,,\,q_{i2}\,$,\,...\}), prompts the target model to answer the questions batch by batch, and records its responses.
In the Tiering stage, JudgeAgent determines the target model’s capability range based on its performance in different batches, thereby deciding the difficulty level on which subsequent question generation in the extension phase should be based. Within each batch, each correct answer makes the total score +1, and the final average score per question is used to evenly divide performance into three tiers: Easy, Medium, and Hard, which then guide adjustments to the question generation rules in the extension phase. In the case that the chosen dataset already provides difficulty labels for its questions (e.g., QuALITY(Pang et al., 2022)), scoring is weighted by difficulty, like Easy +1 and Hard +2, to better reflect the model’s capability level.
@@ 3.2 Interactive Extension
This process constitutes the core of JudgeAgent’s dynamic evaluation, functioning as an “interview” through which JudgeAgent conducts extensive and in-depth evaluations of the target model’s capabilities. Specifically, following the batch divisions determined in the Benchmark Grading stage, JudgeAgent iteratively expands upon each batch by generating additional questions and testing the target model. The extended evaluation proceeds in three steps: Relevant Knowledge Retrieval, Difficulty-Adaptive Question Generation, and Target Testing with Capability Estimation.
@@
Relevant Knowledge Retrieval: Based on the original questions in a batch, JudgeAgent retrieves related knowledge from the dataset’s knowledge base. If only the reference text attached to the original questions is used as background knowledge for generating extended questions, the diversity of generated questions is constrained, leading to high similarity with the original questions and limited coverage of knowledge breadth and depth. Inspired by SoG(Jiang et al., 2025), JudgeAgent also employs a context-graph–based sampling approach for knowledge path retrieval. The context graph is constructed from the dataset’s knowledge base (e.g., the collection of all reference texts) by segmenting texts into chunks and extracting entities as nodes, with entities within the same text range as neighbors. The detailed construction process is provided in Appendix A, and is described in pseudo-code in Algorithm 3.
@@
During knowledge-path sampling, JudgeAgent first extracts entities from the original question as seed entities, then searches the graph for the most similar entity $e$. From the text chunks associated with $e$, JudgeAgent retrieves those most relevant to the question, randomly selects a chunk $c$, and treats $(e,c)$ as the starting node of the path. The entity $e$ is then used as the new seed entity for the next sampling iteration. This process is repeated for $N$ steps to form a knowledge path with both breadth and depth. The text chunks along the path are concatenated to form the background knowledge required for question generation. In this process, the context graph ensures both knowledge breadth and depth, while the greedy similarity strategy in sampling ensures relevance to the original question. Entity extraction is performed by an LLM, with prompts detailed in Appendix C.
@@
Difficulty-Adaptive Question Generation: Unlike other dynamic evaluation methods, JudgeAgent incorporates the target model’s capability estimation into the question-generation process to enable a more fine-grained evaluation. A fixed-difficulty test can serve a screening function by identifying underperforming models, but it offers limited utility in providing a comprehensive evaluation of a model’s knowledge proficiency. For instance, when an overly challenging test is administered to a model with limited knowledge, the evaluation may reveal that the model performs poorly and is indeed below the required competency level, but it fails to delineate the actual range of the target’s knowledge. Therefore, during both the Benchmark Grading and Interactive Extension stages, JudgeAgent dynamically adjusts the estimation of the target’s capability and the difficulty of subsequent questions based on the target’s performance across varying levels of questions. This adaptive approach allows for a precise and refined understanding of the target model’s knowledge boundaries.
@@
Based on the capability estimation obtained from the Benchmark Grading or the previous round of Interactive Extension, JudgeAgent adjusts the difficulty of the generated questions as Easy, Medium, or Hard, and explicitly specifies the generation requirements for each difficulty level in the prompt. For Easy questions, the focus is on assessing information retrieval skills, primarily through cloze-style questions based on the original text. For Medium questions, the emphasis is on understanding concepts and key elements, with questions primarily involving basic inference. For Hard questions, the questions are designed to encourage deep thinking and complex logical analysis. Detailed prompt designs are provided in Prompt 5 in Appendix C.
@@
Target Testing and Capability Estimation: In each round of interactive extension, JudgeAgent presents the generated questions to the target model and collects its responses. Based on responses to all questions from both the Benchmark Grading and Interactive Extension history, JudgeAgent computes scores according to the scoring scheme described in Section 3.1, taking into account both question difficulty and answer correctness. The average score is then used to determine the difficulty level for question generation in the next round of Interactive Extension.

### v2

# JudgeAgent: Knowledge-wise and Dynamic LLM Evaluation with Agent-as-Interviewer

@@ Abstract
Current evaluation paradigms for large language models (LLMs) suffer from overestimated or biased evaluation and mismatched question difficulty, leading to incomplete evaluations of LLM’s knowledge and capability boundaries, which hinder LLM’s effective application and optimization. To address these challenges, we propose Agent-as-Interviewer, a dynamic evaluation paradigm that employs LLM agents to conduct multi-turn interactions for evaluation. Unlike current benchmarking or dynamic interaction paradigms, Agent-as-Interviewer utilizes agents to call knowledge tools for wider and deeper knowledge in the dynamic multi-turn question generation, achieving more complete evaluations of the LLM’s knowledge boundaries. It also leverages agents to plan query strategies for adjustment of the question difficulty levels, enhancing the difficulty control to match the actual capabilities of target LLMs. Based on this paradigm, we develop JudgeAgent, a knowledge-wise dynamic evaluation framework that employs knowledge-driven synthesis as the agent’s tool, and uses difficulty scoring as strategy guidance, thereby finally providing valuable suggestions to help targets optimize themselves. Extensive experiments validate the effectiveness of JudgeAgent’s suggestions, demonstrating that Agent-as-Interviewer can accurately identify the knowledge and capability boundaries of target models. The source code is available on https://anonymous.4open.science/r/JudgeAgent.
@@ 1 Introduction
Evaluating large language models (LLMs) to understand the boundaries of their knowledge and capabilities is a critical step for their successful application across various domains (Tang et al., 2024; Yuan et al., 2023; Shi, 2024; Wang et al., 2025). Current mainstream evaluation methods for LLMs rely on static benchmark evaluations (Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023; Lin et al., 2022; Cobbe et al., 2021; Tang & Yang, 2024), where predefined questions are posed to the target LLM and the target’s performance is evaluated by humans (Chang et al., 2024) or LLM-as-a-judge (Liu et al., 2023; Wang et al., 2023b; Zheng et al., 2023). Benefiting from the controlled question quality and straightforward workflows of the static benchmarking paradigm, developers can rapidly acquire a fundamental understanding of the strengths and limitations of various evolving LLMs, thereby facilitating swift iteration in LLM applications.
However, this benchmarking paradigm faces the bottleneck that benchmarks have become saturated more and more quickly in recent years. It took until 2023 for the accuracy record of MMLU(Hendrycks et al., 2020), which was released in 2020, to reach 80%. While the GPQA benchmark(Rein et al., 2023), published in late 2023, achieved an 80% record just within one year. The bottleneck is due to the static nature(Gu et al., 2024; Ko et al., 2024) of this paradigm. Firstly, the static paradigm restricts evaluations within a predefined knowledge scope, making it difficult to accurately measure the model’s knowledge mastery (Wang et al., 2023a; Kwan et al., 2024), leading to out-of-domain risks in practical applications. Secondly, the lack of dynamic updates increases the risk of data contamination(Schaeffer, 2023; Oren et al., 2023), where target models artificially inflate their benchmark performance by being exposed to test questions during training, leading to an overestimation of their actual capabilities. These shortcomings result in LLMs being misapplied in scenarios beyond their actual capabilities, causing the waste of resources.
Consequently, researchers begin to focus on dynamic evaluation paradigms that leverage LLMs to dynamically modify questions. Initial efforts involve using LLMs to modify questions from static benchmarks to mildly mitigate data contamination (Bai et al., 2023; Shi et al., 2025). Beyond these methods, researchers proposed paradigms that dynamically adjust questions during multi-turn interactions based on the model’s responses, enabling a deeper evaluation (Wang et al., 2023a; Kim et al., 2025). While these dynamic approaches partially address challenges in static paradigms, such as data contamination and plain evaluations, they still suffer from biased evaluation and mismatched difficulty due to simplified feedback mechanisms and knowledge limitations of evaluator LLMs, leading to a misalignment between questions and the actual range of the target’s capability. This misalignment hinders precise evaluations of the target LLM’s knowledge and capability boundaries, thereby impeding effective guidance for both application and subsequent optimization of LLMs.
Figure 1: The difference between Agent-as-Interviewer and current evaluation paradigms.
To address these challenges, we propose Agent-as-Interviewer, a dynamic evaluation paradigm that employs LLM agents to conduct multi-turn interactions for evaluation. Unlike current benchmarking or dynamic interaction paradigms, in the dynamic follow-up question generation based on the target’s responses in multi-turn interactions, Agent-as-Interviewer utilizes agents to call knowledge tools for wider and deeper knowledge, achieving more complete evaluations of the target LLM’s knowledge and capability boundaries. It also leverages agents to plan query strategies for adjustment of the question difficulty levels, enhancing the difficulty control to match the actual capabilities of target LLMs. Based on this paradigm, we develop JudgeAgent, a knowledge-wise dynamic evaluation framework that employs knowledge-driven synthesis as the agent’s tool and uses difficulty scoring as strategy guidance in multi-turn dynamic interactions, thereby finally providing valuable suggestions to help targets optimize themselves.
@@
  • Agent-as-Interviewer paradigm: We propose a dynamic evaluation paradigm that employs LLM agents to conduct multi-turn interactions for evaluation, addressing the challenges of incomplete evaluation and inadequate difficulty control in current paradigms.
  • Knowledge-wise evaluation framework: We introduce JudgeAgent, a knowledge-wise dynamic evaluation framework based on Agent-as-Interviewer, to provide valuable suggestions and precise guidance for the optimization of target LLMs.
  • Validation of the paradigm: We conduct thorough experiments and analysis to validate the effectiveness of Agent-as-Interviewer and JudgeAgent’s suggestions.
@@ 2.1 Static Benchmark-based Evaluation
These methods employ pre-constructed benchmarks to evaluate LLMs in specific tasks, using formats such as multiple-choice, question-answer(Q&A), or prompts for performing tasks. For multiple-choice (Clark et al., 2018; Hendrycks et al., 2021; Huang et al., 2023) or Q&A (Lin et al., 2022; Cobbe et al., 2021; Tang & Yang, 2024) formats, the benchmark provides correct answers and measures the target’s capability by its accuracy. For task-execution format, the benchmark measures the model by the correctness and plausibility assessed by human (Chang et al., 2024) or LLM-as-a-judge (Wang et al., 2023b; Zheng et al., 2023) based on reference answers. These methods ensure controllable question quality and rapid pipelines, but they are also susceptible to data contamination, which undermines the validity of evaluations.
@@ 2.2 Dynamic LLM-based Evaluation
Figure 2: The framework of JudgeAgent. The left part is the interaction process. The central part is the composition of JudgeAgent. The right part presents the tools of JudgeAgent.
These methods leverage LLMs to dynamically generate evaluation questions. Bai et al. (2023) employs LLMs as examiners to generate questions based on Google Trends categories. SafetyQuizzer(Shi et al., 2025) formulates questions with current events retrieved from search engines to maximize the timeliness. These methods focus on the knowledge sources, but lack interaction with target models. Therefore, researchers attempt to generate follow-up questions based on responses in multi-round Q&A sessions. KIEval (Yu et al., 2024) and LLM-as-an-Interviewer (Kim et al., 2025) leverage LLMs to generate follow-up questions and provide feedback based on the target’s responses in multi-turn interactions, thereby probing the depth of the target’s knowledge.
Compared to static methods, their advantage is that target models cannot prepare for generated questions, thereby producing more genuine evaluations. However, these dynamic processes are often centered on a single question, and all the generations rely entirely on the LLM, resulting in a limited scope of knowledge coverage and pool quality control. Moreover, these methods focus on the content of questions, without properly calibrating question difficulty to the target’s capability range, leading to biased evaluations. To address these challenges, we propose Agent-as-Interviewer and develop JudgeAgent based on this paradigm, which performs fine-grained evaluation through knowledge-wise generation and adaptive difficulty adjustment in dynamic multi-turn interactions.
@@ 3 Methodology
We introduce Agent-as-Interviewer, which utilizes LLM agents to simulate the entire interview for the dynamic evaluation of LLMs, and develop JudgeAgent based on this paradigm. In this section, we use JudgeAgent as an example to illustrate the workflow of this paradigm. The workflow comprises three core components: (1) Benchmark Grading: Fundamentally estimating the target’s capability range through testing on public static benchmarks. (2) Interactive Extension: Using knowledge-driven data synthesis to dynamically generate follow-up questions and update the difficulty levels based on the target’s capability estimations. (3) Evaluation Feedback: Evaluating the target’s deficiencies in multiple dimensions and providing actionable suggestions for optimization. The overall framework is illustrated in Figure 2, and detailed as pseudo code in Algorithm 1.
@@ 3.1 Benchmark Grading
Analogous to a written test before an interview, JudgeAgent forms a base understanding of the target model’s capabilities by evaluating its performance on a publicly available static benchmark.
JudgeAgent first partitions the benchmark questions into batches and then prompts the target to answer the questions batch by batch. After assessing the target’s answers, JudgeAgent estimates the target’s capability based on its performance in each batch. Within each batch, each correct answer improves the total score, and the average score determines the follow-up question’s difficulty in the extension stage, including Easy, Medium, and Hard. Let $d_{1},d_{2},d_{3}]$ respectively denote the difficulty Easy, Medium, and Hard, $n_{it},n_{if}$ respectively represent the number of correct and incorrect answers for $d_{i}$-level questions, $c_{it}$ be the score gain for a correct answer at $d_{i}$, and $t_{ij}$ be the score threshold between $d_{i}$ and $d_{j}$. The difficulty control mechanism is formalized as follows:
[…]
In JudgeAgent, we set $c_{1t},c_{2t},c_{3t}]=[1,1.5,2]$ and $t_{12},t_{23}]=[0.5,1]$. For benchmarks without difficulty labels, the default score gain is $c_{2t}$. The mechanism’s detailed design rules, analysis, and proofs are available in Appendix B.
@@ 3.2 Interactive Extension
This process is an ”interview” where JudgeAgent conducts an in-depth evaluation of target models. Specifically, JudgeAgent iteratively generates questions based on the target’s responses and dynamically controls the difficulty. The process consists of three steps: Relevant Knowledge Retrieval, Difficulty-Adaptive Question Generation, and Capability Estimation.
@@
Relevant Knowledge Retrieval: Based on seed question batches, JudgeAgent retrieves related knowledge from the benchmark’s knowledge base. If only reference texts of seed questions are used as background knowledge for generation, the generated questions will be highly similar to the original questions, with limited knowledge scopes. Inspired by SoG(Jiang et al., 2025), JudgeAgent employs a context-graph–based sampling approach to get background texts. The context graph is constructed from the benchmark’s knowledge base (e.g., the collection of all reference texts) by segmenting texts into chunks and extracting entities as nodes, with entities in the same text as neighbors. The detailed construction is illustrated in Appendix C and Algorithm 3.
@@
During sampling, JudgeAgent first extracts entities from seed questions and finds the most similar entity $e$ on the graph. In each hop of sampling, JudgeAgent retrieves the most relevant chunks of $e$ to the question, randomly selects a chunk $c$, treats $(e,c)$ as the root, and $e$ is used as the entity for the next hop. This process is repeated for $N$ hops to form a knowledge path with breadth and depth, and the chunks along the path are concatenated to form the background knowledge. The context graph ensures both knowledge breadth and depth, and the greedy similarity strategy ensures relevance to seed questions. Entity extraction is performed by an LLM, which is GPT-4.1 in our experiments, with prompts detailed in Appendix G. This step provides the appropriate knowledge background for JudgeAgent to generate knowledge-wise questions.
@@
Difficulty-Adaptive Question Generation: Unlike other dynamic methods, JudgeAgent incorporates the target’s capability estimation into question generation to enable more fine-grained evaluations. A fixed-difficulty test can identify underperforming models, but can hardly comprehensively evaluate a model’s knowledge levels. For instance, when an overly challenging test evaluates a model with limited knowledge, the evaluation may reveal that the model performs poorly, but fails to delineate the actual range of its knowledge. Therefore, JudgeAgent dynamically adjusts the target’s capability estimation and the follow-up question’s difficulty based on Eq 1. This adaptive approach allows for a precise and refined understanding of the target’s knowledge boundaries.
@@
Based on the capability estimation, JudgeAgent adjusts the question difficulty as Easy, Medium, or Hard, and explicitly specifies each difficulty’s requirements in the prompt. For Easy, the focus is on assessing information retrieval skills, primarily through cloze-style questions based on the original text, like ”Which virus primarily causes HFMD?”. For Medium, the emphasis is on understanding key concepts, with questions involving basic inference, like ”Which description is correct regarding meiosis II?” For Hard, the questions are designed to encourage deep thinking and complex logical analysis, like ”A patient ingested a toxic substance… The doctor employed… which type of treatment is likely lacking?”. Detailed prompt designs are provided in Appendix G.
@@
Capability Estimation: In each round, JudgeAgent presents the generated questions to the target model and collects its responses. Based on all responses in the above two stages, JudgeAgent computes scores according to the scoring scheme described in Section 3.1, considering both difficulty and correctness. The average score is used to determine the difficulty for the next round’s generation.